# GRADE: Generating multi-hop QA and fine-gRAined Difficulty matrix for RAG Evaluation

### v1

@@ 4.4 Ablation Study
Table 3: Pearson correlation between different aggregation functions of retrieval difficulty $D_{r}$ and answer accuracy, across domains and models.
To verify that the retrieval-difficulty score used in GRADE faithfully tracks task complexity, we ablated the aggregation function that collapses the set of query–chunk embedding similarities into a single scalar. For every question we computed (i) the arithmetic Mean, (ii) a power mean with exponent p=2p=-2 (emphasizing low-similarity entries while retaining signal from the rest), and (iii) the Minimum similarity. We then measured the Pearson correlation between each variant and answer correctness while holding hop depth constant. As Table 3 shows, Min delivers the strongest alignment with accuracy in the Sports domain surpassing the arithmetic mean by roughly 29% and the power mean by 16% on average.
Across the six model–domain pairs, the Min aggregation delivers the highest (or co-highest) Pearson correlation with answer accuracy in four cases and never falls below second place (Table 2). This regularity suggests that overall task difficulty is usually governed by the single hardest retrieval hop: when any required chunk is missed, the reasoning chain cannot be completed and the question is likely answered incorrectly. Because the minimum similarity directly captures this bottleneck, it provides a upper bound on achievable performance.
@@ 5 Conclusion
We present GRADE, a novel evaluation framework for Retrieval-Augmented Generation (RAG) that jointly considers reasoning depth (generator-side) and semantic distance (retriever-side) to better reflect the structural complexity of real-world multi-hop question answering. By defining a difficulty matrix based on hop count and retrieval difficulty, our approach enables fine-grained analysis of RAG systems, disentangling the contributions of retrieval and generation. Experimental results demonstrate that our difficulty metrics strongly correlate with error rates across multiple RAG models in realistic settings. In addition, our knowledge graph augmentation—through missing link detection via semantic clustering—highlights the limitations of traditional triplet-based pipelines in supporting deeper reasoning. Notably, the GRADE difficulty matrix itself serves as a unified difficulty space, where diagonal regions represent scenarios that are jointly challenging for both retrieval and generation. This holistic perspective enables more nuanced diagnoses of RAG system behavior across the full complexity spectrum. Overall, GRADE offers a scalable, interpretable tool for analyzing RAG system failures and guiding future advancements in multi-hop reasoning architectures.
Future work will extend GRADE to enable module-level evaluation across the broader RAG pipeline beyond retrieval and generation, including components such as query decomposition and document filtering.
@@ A.2 Qualitative Results
A.3 Prompts
@@
A.3.1 Prompt for Claim Generation
@@
A.3.2 Prompt for Claim-Sentence Consistency Check
@@
A.3.3 Prompt for Triple Extraction
@@
A.3.4 Prompt for Search Same Entity (Exact / Contextual Equivalence)
@@
A.3.5 Prompt for QA Generation (Ground Truth)
@@ A.3.6 Prompt for Validation
A.3.7 Prompt for Test RAG System Answer Generation

### v2

@@ 4.4 Ablation Study
Table 3: Pearson correlation between different aggregation functions of retrieval difficulty $D_{r}$ and error rate, across domains and models.
To verify that the retrieval-difficulty score used in GRADE faithfully tracks task complexity, we ablated the aggregation function that collapses the set of query–chunk embedding similarities into a single scalar. For every question we computed (i) the arithmetic Mean, (ii) a power mean with exponent p=2p=-2 (emphasizing low-similarity entries while retaining signal from the rest), and (iii) the Minimum similarity. We then measured the Pearson correlation between each variant and error rate while holding hop depth constant. As Table 3 shows, Min delivers the strongest alignment with error rate in the Sports domain surpassing the arithmetic mean by roughly 29% and the power mean by 16% on average.
Across the six model–domain pairs, the Min aggregation delivers the highest (or co-highest) Pearson correlation with error rate in four cases and never falls below second place (Table 2). This regularity suggests that overall task difficulty is usually governed by the single hardest retrieval hop: when any required chunk is missed, the reasoning chain cannot be completed and the question is likely answered incorrectly. Because the minimum similarity directly captures this bottleneck, it provides a upper bound on achievable performance.
@@ 5 Conclusion
We present GRADE, a novel evaluation framework for Retrieval-Augmented Generation (RAG) that jointly considers reasoning depth (generator-side) and semantic distance (retriever-side) to better reflect the structural complexity of real-world multi-hop question answering. By defining a difficulty matrix based on hop count and retrieval difficulty, our approach enables fine-grained analysis of RAG systems, disentangling the contributions of retrieval and generation. Experimental results show that our difficulty metrics strongly correlate with error rates across multiple RAG models in realistic settings. In addition, our knowledge graph augmentation—through missing link detection via semantic clustering—highlights the limitations of traditional triplet-based pipelines in supporting deeper reasoning. Notably, the GRADE difficulty matrix itself serves as a unified difficulty space, where diagonal regions represent scenarios jointly challenging for both retrieval and generation. This holistic perspective enables more nuanced diagnoses of RAG system behavior across the full complexity spectrum. Overall, GRADE offers a scalable, interpretable tool for analyzing RAG system failures and guiding future advancements in multi-hop reasoning architectures.
Future work will extend GRADE to enable module-level evaluation across the broader RAG pipeline, including components such as query decomposition and document filtering.
@@ A.2 Qualitative Results
A.3 Additional Model Families
We extend the evaluation to two non-OpenAI models—Claude 4 Sonnet (20250514) and Llama 3.2 3B Instruct. Both were tested on three representative domains (Sports, Health, Science) under the same setup as described in the main experiment. Across all cases, error rates increase consistently with both hop count and retrieval difficulty ($D_{r}$). This confirms that the proposed difficulty formulation produces stable trends across different model families.
Domain Hops $Q_{1}$ (Low $D_{r}$) $Q_{2}$ $Q_{3}$ $Q_{4}$ (High $D_{r}$)
Sports 2-hop 0.222 0.333 0.241 0.407
3-hop 0.160 0.388 0.592 0.460
4-hop 0.262 0.310 0.429 0.535
5-hop 0.400 0.529 0.600 0.457
Health 2-hop 0.343 0.299 0.522 0.716
3-hop 0.414 0.491 0.534 0.724
4-hop 0.408 0.531 0.633 0.680
5-hop 0.477 0.591 0.659 0.795
Science 2-hop 0.262 0.350 0.525 0.689
3-hop 0.300 0.450 0.517 0.656
4-hop 0.246 0.526 0.456 0.509
5-hop 0.352 0.453 0.648 0.537
Table 5: Error rates for Claude 4 Sonnet (20250514).
Domain Hops $Q_{1}$ (Low $D_{r}$) $Q_{2}$ $Q_{3}$ $Q_{4}$ (High $D_{r}$)
Sports 2-hop 0.519 0.556 0.463 0.630
3-hop 0.420 0.633 0.796 0.720
4-hop 0.452 0.571 0.786 0.860
5-hop 0.514 0.735 0.943 0.686
Health 2-hop 0.262 0.350 0.525 0.689
3-hop 0.300 0.450 0.517 0.656
4-hop 0.246 0.526 0.456 0.509
5-hop 0.352 0.453 0.648 0.537
Science 2-hop 0.508 0.700 0.803 0.836
3-hop 0.533 0.633 0.650 0.738
4-hop 0.561 0.702 0.649 0.702
5-hop 0.537 0.698 0.815 0.685
Table 6: Error rates for Llama 3.2 3B Instruct.
A.4 Evaluation on Multi-Step and Decomposition RAG Systems
We extend the evaluation to two stronger RAG pipelines under the same protocol and difficulty matrix as in the main paper. (1) Multi-step RAG: two-stage retrieval and generation—first generate an intermediate answer from top-10 retrieved documents, then perform a second retrieval conditioned on the original query, retrieved context, and the intermediate answer to produce the final response. (2) Query-decomposition RAG: decompose the query into sub-queries; for each sub-query retrieve top-10 documents, merge with de-duplication, and rerank using bge-reranker-large. All results are reported as cell-wise error rates $(1-\text{accuracy})$ over the $\times4$ difficulty matrix (rows = hop count; columns = $D_{r}$ quartiles from low to high).
In both systems and across domains, error rates generally increase with hop count and retrieval difficulty, with the largest gains along the main diagonal of the matrix (higher hops and higher $D_{r}$).
A.5 Evaluation on Entertainment and Technology Domains
To further examine the robustness of our framework across domains, we extended the experiments to Entertainment and Technology corpora. Following the same pipeline described in Section 3, we constructed domain-specific knowledge graphs and generated multi-hop queries (2–5 hops). For both domains, error rates showed consistent trends with our main analysis: performance degrades as reasoning depth increases and retrieval difficulty grows, while overall tendencies align with the Sports, Health, and Science results.
As shown in Table 9 and Table 10, these additional experiments reinforce the generalizability of the proposed difficulty framework across diverse topical domains. The observed patterns mirror those in the primary analysis, further supporting the consistency of GRADE under varying domain conditions.
A.6 Human Evaluation on Missing Link Detection
To assess the reliability of our missing link detection process, we manually analyzed 30 randomly sampled entity groups from the augmentation phase. We observed 5 ambiguous cases among exact-equivalence groups and 6 among contextually-equivalent groups. Most were not outright errors but borderline situations involving partial overlaps or differences in granularity. For example, “the first round of the draft” versus “the first round of the NFL draft” reflected an inclusion rather than strict synonymy, while “redemption” versus “Michael Vick’s ongoing redemption” showed a mismatch in abstraction level. Such cases were limited in scope relative to the overall mappings and did not noticeably affect question generation or evaluation outcomes. Still, we note that refining entity matching strategies remains a worthwhile direction, especially in high-stakes domains.
Multi-step RAG Hops $Q_{1}$ (Low $D_{r}$) $Q_{2}$ $Q_{3}$ $Q_{4}$ (High $D_{r}$)
Sports 2-hop 0.352 0.296 0.185 0.444
3-hop 0.180 0.388 0.490 0.420
4-hop 0.357 0.333 0.476 0.465
5-hop 0.371 0.706 0.600 0.457
Health 2-hop 0.403 0.299 0.507 0.672
3-hop 0.310 0.421 0.466 0.569
4-hop 0.388 0.490 0.592 0.660
5-hop 0.591 0.500 0.568 0.727
Science 2-hop 0.262 0.383 0.557 0.721
3-hop 0.367 0.483 0.550 0.705
4-hop 0.316 0.491 0.439 0.509
5-hop 0.352 0.472 0.630 0.537
Table 7: Extended evaluation on a two-stage Multi-step RAG. Values are error rates over the $\times4$ difficulty matrix (rows = hop count; columns = $D_{r}$ quartiles). Lower is better.
Query-decomposition RAG Hops $Q_{1}$ (Low $D_{r}$) $Q_{2}$ $Q_{3}$ $Q_{4}$ (High $D_{r}$)
Sports 2-hop 0.519 0.500 0.426 0.500
3-hop 0.360 0.531 0.653 0.580
4-hop 0.500 0.738 0.738 0.605
5-hop 0.514 0.588 0.629 0.743
Health 2-hop 0.478 0.493 0.627 0.746
3-hop 0.690 0.754 0.776 0.672
4-hop 0.408 0.714 0.714 0.740
5-hop 0.705 0.750 0.750 0.773
Science 2-hop 0.443 0.500 0.623 0.787
3-hop 0.550 0.533 0.717 0.705
4-hop 0.491 0.684 0.579 0.649
5-hop 0.648 0.528 0.630 0.741
Table 8: Extended evaluation on a Query-decomposition RAG with per-subquery retrieval, merge & de-duplication, and bge-reranker-large reranking. Values are error rates.
Hops Low $D_{r}$ Q2 Q3 High $D_{r}$
2-hop 0.375 0.298 0.489 0.479
3-hop 0.471 0.440 0.549 0.608
4-hop 0.378 0.523 0.689 0.578
5-hop 0.708 0.522 0.652 0.542
Table 9: Error rates in the Entertainment domain. Lower is better.
Hops Low $D_{r}$ Q2 Q3 High $D_{r}$
2-hop 0.431 0.380 0.420 0.608
3-hop 0.417 0.375 0.542 0.667
4-hop 0.386 0.465 0.488 0.545
5-hop 0.564 0.564 0.487 0.550
Table 10: Error rates in the Technology domain. Lower is better.
A.7 Evaluation with Alternative Embedding Models
Hops Low $D_{r}$ Q2 Q3 High $D_{r}$
all-mpnet-base-v2
2-hop 0.296 0.121 0.241 0.389
3-hop 0.220 0.388 0.469 0.360
4-hop 0.262 0.262 0.452 0.442
5-hop 0.286 0.441 0.571 0.514
bge-en-large-v1.5
2-hop 0.315 0.111 0.222 0.333
3-hop 0.180 0.286 0.449 0.400
4-hop 0.191 0.310 0.405 0.442
5-hop 0.371 0.265 0.429 0.686
jina-embeddings-v3
2-hop 0.315 0.167 0.259 0.444
3-hop 0.140 0.388 0.469 0.600
4-hop 0.167 0.357 0.595 0.651
5-hop 0.400 0.353 0.543 0.771
Table 11: Error rates in the Sports domain with three embedding models.
Hops Low $D_{r}$ Q2 Q3 High $D_{r}$
all-mpnet-base-v2
2-hop 0.343 0.254 0.388 0.731
3-hop 0.310 0.421 0.552 0.603
4-hop 0.347 0.531 0.510 0.600
5-hop 0.386 0.659 0.477 0.545
bge-en-large-v1.5
2-hop 0.299 0.299 0.388 0.612
3-hop 0.362 0.404 0.483 0.448
4-hop 0.347 0.531 0.633 0.460
5-hop 0.500 0.523 0.682 0.477
jina-embeddings-v3
2-hop 0.313 0.269 0.463 0.716
3-hop 0.345 0.404 0.431 0.655
4-hop 0.306 0.571 0.531 0.560
5-hop 0.432 0.682 0.568 0.636
Table 12: Error rates in the Health domain for three emedding models.
Hops Low $D_{r}$ Q2 Q3 High $D_{r}$
all-mpnet-base-v2
2-hop 0.230 0.367 0.492 0.689
3-hop 0.383 0.383 0.633 0.574
4-hop 0.228 0.474 0.456 0.544
5-hop 0.333 0.547 0.519 0.537
bge-en-large-v1.5
2-hop 0.295 0.367 0.393 0.705
3-hop 0.333 0.500 0.400 0.623
4-hop 0.263 0.386 0.404 0.404
5-hop 0.352 0.358 0.500 0.481
jina-embeddings-v3
2-hop 0.213 0.333 0.443 0.607
3-hop 0.367 0.500 0.433 0.639
4-hop 0.246 0.404 0.439 0.544
5-hop 0.296 0.453 0.481 0.611
Table 13: Error rates in the Science domain for three embedding models.
To further assess the robustness of the GRADE framework, we conducted additional experiments using three widely adopted embedding encoders: all-mpnet-base-v2  [^4] https://huggingface.co/sentence-transformers/all-mpnet-base-v2, bge-en-large-v1.5  [^5] https://huggingface.co/BAAI/bge-large-en-v1.5, and jina-embeddings-v3  [^6] https://huggingface.co/jinaai/jina-embeddings-v3. These models differ substantially in architecture, training data, and design philosophy, providing complementary perspectives on sentence-level representation.
As shown in Tables 11, Tables 12 and Tables 13, Across all domains (Sports, Health, Science), we observed consistent trends aligned with our main findings: diagonal values in the GRADE matrix exhibit a steady increase from upper-left to bottom-right. This consistency across encoders demonstrates that the framework’s evaluation signals are not artifacts of a particular embedding choice, but rather reflect intrinsic aspects of multi-hop reasoning difficulty.
A.8 Evaluation under Alternative Chunking Strategies
To examine whether the GRADE framework is sensitive to specific retrieval granularities, we evaluated two alternative chunking strategies in addition to the main configuration: 64--128 tokens, 25-token overlap and 256--512 tokens, 100-token overlap.
As shown in Table 14, Table 15 and Table 16, across all domains (Sports, Health, Science), we observed the same characteristic patterns as in our main experiments: diagonal values in the GRADE matrix exhibit a steady increase from upper-left to bottom-right. This indicates that the validity of the framework’s difficulty signals does not depend on a particular chunking configuration.
Hops Low $D_{r}$ Q2 Q3 High $D_{r}$
256--512 tokens, 100-token overlap
2-hop 0.278 0.315 0.204 0.407
3-hop 0.140 0.388 0.673 0.340
4-hop 0.190 0.286 0.381 0.372
5-hop 0.314 0.529 0.543 0.429
64--128 tokens, 25-token overlap
2-hop 0.222 0.259 0.204 0.333
3-hop 0.120 0.490 0.673 0.400
4-hop 0.286 0.333 0.619 0.488
5-hop 0.429 0.529 0.714 0.486
Table 14: Error rates in the Sports domain with alternative chunking strategies.
Hops Low $D_{r}$ Q2 Q3 High $D_{r}$
256--512 tokens, 100-token overlap
2-hop 0.373 0.224 0.433 0.672
3-hop 0.310 0.351 0.466 0.603
4-hop 0.327 0.429 0.490 0.580
5-hop 0.455 0.455 0.432 0.682
64--128 tokens, 25-token overlap
2-hop 0.403 0.239 0.537 0.642
3-hop 0.310 0.386 0.517 0.586
4-hop 0.367 0.531 0.531 0.660
5-hop 0.455 0.523 0.500 0.727
Table 15: Error rates in the Health domain with alternative chunking strategies.
Hops Low $D_{r}$ Q2 Q3 High $D_{r}$
256--512 tokens, 100-token overlap
2-hop 0.246 0.333 0.475 0.689
3-hop 0.350 0.417 0.483 0.672
4-hop 0.228 0.491 0.456 0.526
5-hop 0.352 0.415 0.519 0.574
64--128 tokens, 25-token overlap
2-hop 0.213 0.350 0.475 0.672
3-hop 0.383 0.400 0.567 0.525
4-hop 0.246 0.456 0.368 0.544
5-hop 0.315 0.547 0.537 0.574
Table 16: Error rates in the Science domain with alternative chunking strategies.
A.9 Prompts
@@
A.9.1 Prompt for Claim Generation
@@
A.9.2 Prompt for Claim-Sentence Consistency Check
@@
A.9.3 Prompt for Triple Extraction
@@
A.9.4 Prompt for Search Same Entity (Exact / Contextual Equivalence)
@@
A.9.5 Prompt for QA Generation (Ground Truth)
@@ A.9.6 Prompt for Validation
A.9.7 Prompt for Test RAG System Answer Generation